OpenStudy (anonymous):

Select the ordered pair from the choices below that is a solution to the following system of equations: 3x = y + 18 y + 7 = -30x (7, 3) (-2, 12) (1/3, -17) (-4, 2)

OpenStudy (anonymous):

Answer:(1/3, -17) Subtract the two equations 3x=y+18 -(-30x=y+7) ----------- 33x=11 x=11/33 x=1/3 Plug in to get Y: 3*(1/3)=y+18 1=y+18 y=-17 So, (1/3, -17)